OpenStudy (anonymous):

square root to the 3rd power n^2 -1+4 =3

OpenStudy (anonymous):

\[\sqrt[3]{n ^{2}-1}+4=3?\]

OpenStudy (anonymous):

\[\sqrt[3]{n^2 -1+4}=3\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

cool. Cube both sides first.

OpenStudy (anonymous):

Wait. Which equation? Yours or mine?

OpenStudy (anonymous):

they're the same.

OpenStudy (anonymous):

could give me a visual of cubing both sides.

OpenStudy (anonymous):

Mine has the +4 on the outside of the root. Yours has it on the inside. Which one is it?

OpenStudy (anonymous):

If it's on the outside, then you'll subtract 4 first. If it's on the inside, then you'll cube both sides first.

OpenStudy (anonymous):

yours

OpenStudy (anonymous):

Okay. So subtract 4 first and you'll get...\[\sqrt[3]{n ^{2}-1}=-1\]

OpenStudy (anonymous):

Then cube both sides and you'll get...\[n ^{2}-1=-1\]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

Next, add 1 to both sides and you'll get...\[n ^{2}=0\]

OpenStudy (anonymous):

Therefore, n = 0.

OpenStudy (anonymous):

Which if you plug it into the original equation, it works. You can take the cube root of negative numbers, but you can't take the square root. As a matter of fact: You can take the ODD-number root of any negative number. But you CANNOT take the EVEN-number root of any negative number. Just a rule to keep in mind. Take care.

OpenStudy (anonymous):

thank you.